OpenStudy (anonymous):

-5 - 3 = -2 rightt ?

OpenStudy (anonymous):

No thats wrong, its -8.

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

raheen whats wrong??

OpenStudy (anonymous):

I kneww itt, my teacher said that was wrong is the only part of the equation she said I needed to check @

OpenStudy (anonymous):

oh, read in deffernt way, it's -8 not -2

myininaya (myininaya):

-5-3=-8

myininaya (myininaya):

\[5+3=8 => -(5+3)=-8 \]

OpenStudy (anonymous):

yeaahh I thought it was -8 but she told me it was wrong so I tried -2

myininaya (myininaya):

\[-5+3=-2\]

myininaya (myininaya):